OpenStudy (therealmeeeee):

Insert grouping symbols so that the given expression has a value of 22. 10 – 5 • 8 – 6 + 12 A. (10 – 5 • 8) – 6 + 12 B. (10 – 5) • 8 – (6 + 12) C. 10 – 5 • (8 – 6) + 12 D. (10 – 5) • (8 – 6 + 12)

OpenStudy (therealmeeeee):

@AmandaMichelle

OpenStudy (therealmeeeee):

anyone going to help me????

OpenStudy (anonymous):

10 - 5 * 8 - 6 + 12 = 22 (10 - 5) * (8 - 6) + 12 = 22 5 * 2 + 12 = 22 10 + 12 = 22 22 = 22